13 January 2014 – Monday

LDAP:

  • Issues from over the weekend:
    • Friday: An ID card barcode seems to be owned by two accounts.  This usually means that we have found a person with two LDAP entries (i.e. we have had a failure in the IDM process to properly match to roles for the same person).  However, this time, the problem is that we had a duplication that we found and corrected, and the CatCard process has not gotten that fact merged yet.  Notified the CatCard office of the issue.

Backups:

  • Fulls from the weekend that are still running:  NFS Fileserver 2, B-H Library Full, COMIS MED23 A.
    • The B-H Library Full is a surprise.  Seems we have a filesystem that by name makes me think it is related to an Oracle database that has been backing up for over 50 hours and has managed to send 1276GB so far.  So, I emailed the people that I think are the system administrators to check up on if this is expected and if it is truly an Oracle database (in which case this is not the way to properly back it up).
  • Last week’s DISK clones are still going.  They have 21,974.153 GB left to write.  I’ll let that chug along until the NFS2 backup finally completes – then kill it and get it restarted to pull in this week’s clones as well.
  • Ran the TAPE clones – they completed in a couple of hours (not much happens in the tape world on the second weekend).
  • Moving tapes around from the library to the vaults and back.  Sixty-four (64) tapes taken to the vaults.  One hundred twenty (120) tapes brought back for re-use.

Tags:

Leave a Reply

You must be logged in to post a comment.